Silicon has three common isotopes. The masses of the first two isotopes and percent abundances are as follows: 27.977 amu, 92.223% and 28.976 amu, 4.685%.

If the atomic mass of silicon is 28.085 amu, what is the percent abundance of the third isotope? What is its atomic mass?
1
Expert's answer
2020-11-04T14:24:01-0500

W=100-92.223-4.685=3.092%

27.977×0.92223+28.976×0.04685+0.03092•x=28.085,

0.03092•x=28.085-25.80123-1.35753,

0.03092•x=0.92624,

x=29.956